OpenStudy (anonymous):

How to solve 2X^2+4X+4=0 by completing the square. Please explain thoroughly!

hero (hero):

2x^2 + 4x + 4 = 0 2(x^2 + 2x + 2) = 0 x^2 + 2x + 2 = 0 x^2 + 2x = -2 x^2 + 2x + 1 = -2 + 1 (x+1)^2 = -1 x + 1 = + or - sqrt{-1} x = -1 + i x = -1 - i

OpenStudy (anonymous):

Ah perfect! I was wondering what to do with the factor and whether or not it was negative. A well deserved medal mate!

OpenStudy (anonymous):

Wait does that say I hate math...

hero (hero):

Just an avatar mate

OpenStudy (anonymous):

Meh I can't say I am in love with mah truthfully. Heck I think math hates me as much as I hate it! ;)

hero (hero):

Okay, good luck bro.

OpenStudy (anonymous):

Thanks for your help!